Please help me solve this word problem:
Glenna wants to rent a car for a trip to Key West for one week. She calls two car rental companies to get prices. Mr. Kotter's Rentals rents a Cadillac Escalade for $99 foe one week plus $0.11 per mile over 100 miles. Barbino's Rentals has the same vehicle for $75 per week and $0.15 per mile over 150 miles. If it is 432 miles to Key West, which company has the better deal? How many miles would Glenna drive before she would be spending the same amount at either company?
----Thank you =)

Answer by ankor@dixie-net.com(6693)

Glenna wants to rent a car for a trip to Key West for one week.
She calls two car rental companies to get prices.
Mr. Kotter's Rentals rents a Cadillac Escalade for $99 foe one week plus $0.11 per mile over 100 miles.
Barbino's Rentals has the same vehicle for $75 per week and $0.15 per mile over 150 miles.
If it is 432 miles to Key West, which company has the better deal?
:
Find the cost of each for a 432 mi trip
K = .11(432-100) + 99
K = .11(332) + 99
K = 36.52 + 99
K = $135.52
:
B = .15(432-150) + 75
B = .15(282) + 75
B = 42.30
B = $117.30
:
How many miles would Glenna drive before she would be spending the same amount
at either company?
Let m = miles
K's cost = B's cost
.11(m-100) + 99 = .15(m-150) + 75
.11m - 11 + 99 = .15m - 22.50 + 75
.11m + 88 = .15m + 52.50
88 - 52.50 = .15m - .11m
35.5 = .04m
m =
m = 887.5 miles to cost the same